OpenStudy (anonymous):

If 27.31mL of 0.2115 M NaOH is able to nuetralize 37.45mL of HCI what is the concentration of the acid?

OpenStudy (anonymous):

@jovis2009 In order for neutralization to occur there must be equal number of normals of acid and base. Since the equivalent weight is the same as the molecular weights for NaOH and HCl the normality and the molarities are the same. Normality of NaOH =N1=0.2115 N Volume of NaOH =V1=27.31 ml Normality of HCl =N2=x N Volume of HC l=37.45 ml. So according to the above stated law, V1*N1=V2*N2 27.31*0.2115=37.45*x , x=0.1542 N=0.1542 M . Therefore the concentration of HCl is 0.1542 moles/liter.

OpenStudy (anonymous):

why didnt you change mL to L?

OpenStudy (anonymous):

Molarity is a unit which measures concentration in moles/liter so moles/liter and M are the same units I didnt change anything.

OpenStudy (anonymous):

No need to change it to liters ml works fine in the above formula.

OpenStudy (anonymous):

ok thank u very much

OpenStudy (anonymous):

Yw.If you want to convert the volume to liters, you have to divide both V1 and V2 by 1000.If you notice the equation they will cancel out.

OpenStudy (anonymous):

ok then how will you find volume of 0.11 M HCI needed to neutralize 28.67 of 0.137M KOH?

OpenStudy (anonymous):

35.707 ml is required.The same formula applies here.